# Inducing Programmatic Skills for Agentic Tasks

### v1

@@ 2.3 Inducing and Verifying Programmatic Skills
After rewriting and truncation, we get this skill-using trajectory $\tau_{D}$ as the prefix to test skills. We now query the agent $\pi$ again with the same NL query $q$ and first execute $\tau_{D}$ on the environment. We then allow agents to continue generating up to $H_{max}-|\tau_{D}|$ actions to finish the task. In the Figure 2 example, to successfully solve query $q$, we expect the agent to generate another step of send_msg_to_user(...) with the correct answer ‘2’ in the message. We take the concatenation of the trajectory prefix $\tau_{D}$ and the later additionally produced new steps $\tau_{A}$ (e.g., [[[[send_msg_to_user(‘2’)]]]]) as the full trajectory $\tau_{f}$. We then decide whether to add the induced programs $\mathcal{D}$ into the agent skill library as $\mathcal{A}$ by examining $\tau_{f}$.
[…]
@@ 3.2 Results and Analysis
As shown in Table 1, compared to the vanilla static-agent baseline, adaptive agents (AWM and ASI) generally achieve 11.0–23.5% higher success rates overall. Among adaptive agents, our ASI with programmatic skills, achieves another substantial success rate gain of 11.3% across websites, compared to its AWM counterpart that induces and uses textual skills. Meanwhile, ASI offers additional efficiency benefits by reducing the number of steps in solutions by 15.3% and 10.6% than vanilla and AWM agents, as one skill-call action can often execute multiple steps written in primitive actions used by vanilla and AWM agents. These advantages in correctness and efficiency are exhibited prominently across different websites and tasks, as shown by the website breakdown on Table 1 (right).
@@ 4 Scaled-Up Browsing Activities
Therefore, we curate several case scenarios where the user asks for action-dense instructions, such as the tasks listed in Figure 3. Because the tasks are long-horizon and involve multiple sub-tasks, we follow Xu et al. (2024) and set up intermediate checkpoints to better track the intermediate progress of agents. Refer to §B.1 to see the full list of tasks and their evaluation checkpoints. We measure the success rate of each example by the percentage of checkpoints achieved by the agent. We report the average success rate of all examples, as well as the average number of steps taken to solve the tasks, in Table 4.
@@
Table 4: Performance of vanilla, AWM, and ASI agents in scaled-up browsing scenarios. We perform statistical testing between ASI and each baseline and verify all improvements are statistically significant with t-statistics $|t|>2$ and $p<0.05$; see §B.3 for more details.
@@ 5 Adapting Across Websites
For each sandbox-real website pair, we take ten information-seeking style queries (He et al., 2024) in WebArena that do not involve potential privacy leakage or unrecoverable risky actions, such as making a purchase or changing user password. We provide the task details in §B.2. We compare ASI and AWM with their programmatic and textual skills as learned in §3, as well as comparing to the vanilla static agent baseline.
@@
Table 6: Cross-website results. ASI significantly surpasses baselines in sr and # steps (with $|t|>2$ and $p<0.05$) from our analysis in §B.3.
@@ A.2 Prompts for LLM-Based Components
Prompt for Episode Evaluation You are an expert in evaluating the performance of a web navigation agent. The agent is designed to help a human user navigate a website to complete a task. Given the user’s intent, the agent’s action history, the final state of the webpage, and the agent’s response to the user, your goal is to decide whether the agent’s execution is successful or not. Please be careful of each detail and strict about the evaluation process. There are three types of tasks: 1. Information seeking: The user wants to obtain certain information from the webpage, such as the information of a product, reviews, map info, comparison of map routes, etc. The bot’s response must contain the information the user wants, or explicitly state that the information is not available. Otherwise, e.g. the bot encounters an exception and respond with the error content, the task is considered a failure. Besides, be careful about the sufficiency of the agent’s actions. For example, when asked to list the top-searched items in a shop, the agent should order the items by the number of searches, and then return the top items. If the ordering action is missing, the task is likely to fail. 2. Site navigation: The user wants to navigate to a specific page. Carefully examine the bot’s action history and the final state of the webpage to determine whether the bot successfully completes the task. No need to consider the bot’s response. 3. Content modification: The user wants to modify the content of a webpage or configuration. Carefully examine the bot’s action history and the final state of the webpage to determine whether the bot successfully completes the task. No need to consider the bot’s response. *IMPORTANT* Please be strict about the evaluation process. Format your response into two lines as shown below: Thoughts: $<$your thoughts and reasoning process$>$ Status: ”success” or ”failure”
@@
Appendix B Scaled-Up and Cross-Website Tests
We provide the full list of tasks used in scaled-up (§4) and cross-website (§5) analyses in §B.1 and §B.2, respectively. In §B.3, we further perform significance testing to validate the findings from Table 4 and Table 6.
B.1 Scaled-Up Tasks
@@
B.2 Cross-Website Tasks

### v2

@@ 2.3 Inducing and Verifying Programmatic Skills
After rewriting and truncation, we get this skill-using trajectory $\tau_{D}$ as the prefix to test skills. We now query the agent $\pi$ again with the same NL query $q$ and first execute $\tau_{D}$ on the environment. We then allow agents to continue generating up to $H_{max}-|\tau_{D}|$ actions to finish the task. In the Figure 2 example, to successfully solve query $q$, we expect the agent to generate another step of send_msg_to_user(...) with the correct answer ‘2’ in the message. We take the concatenation of the trajectory prefix $\tau_{D}$ and the later additionally produced new steps $\tau_{A}$ (e.g., [[send_msg_to_user(‘2’)]]) as the full trajectory $\tau_{f}$. We then decide whether to add the induced programs $\mathcal{D}$ into the agent skill library as $\mathcal{A}$ by examining $\tau_{f}$.
Specifically, we check $\tau_{f}$ from three dimensions: (1) Correctness: if executing $\tau_{f}$ successfully solves the task $q$ as judged by the neural model evaluator $V_{\mathcal{L}}$; (2) Skill Usage: if the trajectory contains at least one call to at least one new skill in $\mathcal{D}$; and (3) Skill Validity: if all skill-calling actions cause environment changes. If all three boxes are checked, we add the skills being called in the trajectory $\tau_{f}$ to the agent skill library $\mathcal{A}_{t}\cup\mathcal{D}_{\it called}\rightarrow\mathcal{A}_{t+1}$. By adding $\mathcal{D}_{\it called}$, the agent can now generate actions that call these skill programs to solve subsequent tasks.
@@ 3.2 Results and Analysis
In Table 1, compared to the vanilla static-agent baseline, adaptive agents (AWM and ASI) generally achieve 11.0–23.5% higher success rates overall. Among adaptive agents, our ASI with programmatic skills, achieves another 11.3% success rate gain across websites, compared to its AWM counterpart that induces and uses textual skills. Meanwhile, ASI offers additional efficiency benefits by reducing the number of steps in solutions by 15.3% and 10.6% than vanilla and AWM agents, as one skill-call action can often execute multiple steps written in primitive actions used by vanilla and AWM agents. These advantages in correctness and efficiency are exhibited prominently across different websites and tasks, as shown by the website breakdown on Table 1 (right). Refer to §B for more analysis.
@@ 4 Scaled-Up Browsing Activities
Therefore, we curate several case scenarios where the user asks for action-dense instructions, such as the tasks listed in Figure 3. Because the tasks are long-horizon and involve multiple sub-tasks, we follow Xu et al. (2024) and set up intermediate checkpoints to better track the intermediate progress of agents. Refer to §C.1 to see the full list of tasks and their evaluation checkpoints. We measure the success rate of each example by the percentage of checkpoints achieved by the agent. We report the average success rate of all examples, as well as the average number of steps taken to solve the tasks, in Table 4.
@@
Table 4: Performance of vanilla, AWM, and ASI agents in scaled-up browsing scenarios. We perform statistical testing between ASI and each baseline and verify all improvements are statistically significant with t-statistics $|t|>2$ and $p<0.05$; see §C.3 for more details.
@@ 5 Adapting Across Websites
For each sandbox-real website pair, we take ten information-seeking style queries (He et al., 2024) in WebArena that do not involve potential privacy leakage or unrecoverable risky actions, such as making a purchase or changing user password. We provide the task details in §C.2. We compare ASI and AWM with their programmatic and textual skills as learned in §3, as well as comparing to the vanilla static agent baseline.
@@
Table 6: Cross-website results. ASI significantly surpasses baselines in sr and # steps (with $|t|>2$ and $p<0.05$) from our analysis in §C.3.
@@ A.2 Prompts for LLM-Based Components
Prompt for Episode Evaluation You are an expert in evaluating the performance of a web navigation agent. The agent is designed to help a human user navigate a website to complete a task. Given the user’s intent, the agent’s action history, the final state of the webpage, and the agent’s response to the user, your goal is to decide whether the agent’s execution is successful or not. Please be careful of each detail and strict about the evaluation process. There are three types of tasks: 1. Information seeking: The user wants to obtain certain information from the webpage, such as the information of a product, reviews, map info, comparison of map routes, etc. The bot’s response must contain the information the user wants, or explicitly state that the information is not available. Otherwise, e.g. the bot encounters an exception and respond with the error content, the task is considered a failure. Besides, be careful about the sufficiency of the agent’s actions. For example, when asked to list the top-searched items in a shop, the agent should order the items by the number of searches, and then return the top items. If the ordering action is missing, the task is likely to fail. 2. Site navigation: The user wants to navigate to a specific page. Carefully examine the bot’s action history and the final state of the webpage to determine whether the bot successfully completes the task. No need to consider the bot’s response. 3. Content modification: The user wants to modify the content of a webpage or configuration. Carefully examine the bot’s action history and the final state of the webpage to determine whether the bot successfully completes the task. No need to consider the bot’s response. *IMPORTANT* Please be strict about the evaluation process. Format your response into two lines as shown below: Thoughts: <<your thoughts and reasoning process>> Status: ”success” or ”failure”
@@
Appendix B Skill Induction: Analysis
We provide more details about the skill induction process, in skill curation and reuse frequency (§B.1) and representative skill case studies (§B.2)
B.1 Skill Induction and Reusability
To provide more insights on how agents curate and reuse programmatic skills, for the main experiments on WebArena, we calculate the number of examples that (i) attempt to induce a new skill, (ii) successfully induce a new skill, and (iii) reuse a previously induced skill.
As shown in Table 8, the agent reuses at least 1 skill for 42.5% of the cases, which is pretty frequent across websites. Moreover, these skills are created using 7.6% of the tasks, demonstrating the high scalability of our skill learning approach.
Domain Number of Examples
Attempted Induction Successful Induction Reuse Examples Total
shopping 21 8 96 180
admin 38 15 108 156
reddit 24 11 14 105
map 13 10 80 109
gitlab 25 11 10 175
Table 8: Analyzing the number of examples that attempt, induce, and reuse skills.
B.2 Representative Skills
We demonstrate two representative types of skills, which (i) chain frequently-used-together actions, and (ii) involve advanced programming primitives.
Chaining Primitive Actions
The most common type of skills chains multiple frequently-used-together primitive actions (e.g, click, fill) in a certain order, to reflect a particular common usage, such as the search_product tool illustrated below. This type of skill provides an easy portal for frequent workflows.
1def search_product(name: str):
2 click("Search textbox...")
3 if has_popup_window():
4 click("Close")
5 fill("Search textbox", name)
6 click("Search button")
Advanced Programming Primitives
Besides a sequential chaining of actions, some skills contain more advanced logics, such as if-else or for/while loops, as the two examples below.
1def navigate_and_sort_category(category_id: str, subcategory_id: str | None = None):
2 """Navigate to a product category and sort items by price.
3
4 Args:
5 category_id: ID of the main category link
6 subcategory_id: Optional ID of the subcategory link, if needed
7
8 Returns:
9 None
10
11 Examples:
12 navigate_and_sort_category(’1241’, ’1873’) # PS4 under Video Games
13 navigate_and_sort_category(’1245’) # For main category only
14 """
15 click(category_id) # Click main category
16 if subcategory_id:
17 click(subcategory_id) # Click subcategory if provided
18 select_option("1553", "Price") # Sort by price ascending
1def browse_category_by_navigation(menu_id_sequence: list):
2 """Browse products by navigating through a sequence of menu IDs.
3 This function allows navigation through a series of menu interactions.
4 Args:
5 menu_id_sequence (list): A list of menu IDs to interact sequentially, using hover actions followed by a click.
6
7
8 Example usage:
9 browse_category_by_navigation([’735’, ’786’, ’797’]) # Navigates Home & Kitchen -> Storage & Organization -> Racks, Shelves & Drawers
10 """
11 for idx, menu_id in enumerate(menu_id_sequence[:-1]):
12 hover(menu_id)
13 click(menu_id_sequence[-1]) # Click the final id to land in the predefined category
Appendix C Scaled-Up and Cross-Website Tests
We provide the full list of tasks used in scaled-up (§4) and cross-website (§5) analyses in §C.1 and §C.2, respectively. In §C.3, we further perform significance testing to validate the findings from Table 4 and Table 6.
C.1 Scaled-Up Tasks
@@
C.2 Cross-Website Tasks